# An Adaptive Online Smoother with Closed-Form Solutions and Information-Theoretic Lag Selection for Conditional Gaussian Nonlinear Systems

### v1

@@ 2.1 Conditionally Gaussian Processes
In its most general form, a conditional Gaussian system of processes consists of two diffusion-type processes defined by the following system of stochastic differentials given in Itô form [36, 38, 50]:
$\mathrm{d}\mathbf{x}(t)$ $=\left[\boldsymbol{\Lambda}^{\mathbf{x}}(t,\mathbf{x})\mathbf{y}(t)+\mathbf{f}^{\mathbf{x}}(t,\mathbf{x})\right]\mathrm{d}t+\boldsymbol{\Sigma}_{1}^{\mathbf{x}}(t,\mathbf{x})\mathrm{d}\mathbf{W}_{1}(t)+\boldsymbol{\Sigma}_{2}^{\mathbf{x}}(t,\mathbf{x})\mathrm{d}\mathbf{W}_{2}(t),$ (2.1)
$\mathrm{d}\mathbf{y}(t)$ $=\left[\boldsymbol{\Lambda}^{\mathbf{y}}(t,\mathbf{x})\mathbf{y}(t)+\mathbf{f}^{\mathbf{y}}(t,\mathbf{x})\right]\mathrm{d}t+\boldsymbol{\Sigma}_{1}^{\mathbf{y}}(t,\mathbf{x})\mathrm{d}\mathbf{W}_{1}(t)+\boldsymbol{\Sigma}_{2}^{\mathbf{y}}(t,\mathbf{x})\mathrm{d}\mathbf{W}_{2}(t),$ (2.2)
@@
are two mutually independent complex-valued Wiener processes (i.e., both their real and imaginary parts are mutually independent real-valued Wiener processes) and almost every path of $\mathbf{x}$ and $\mathbf{y}$ is in $C^{0}([0,T];\mathbb{C}^{k})$ and $C^{0}([0,T];\mathbb{C}^{l})$, respectively. The elements of the vector- and matrix-valued functions of multiplicative factors ($\boldsymbol{\Lambda}^{\mathbf{x}},\boldsymbol{\Lambda}^{\mathbf{y}}$), forcings ($\mathbf{f}^{\mathbf{x}},\mathbf{f}^{\mathbf{y}}$), and noise feedbacks ($\boldsymbol{\Sigma}_{1}^{\mathbf{x}},\boldsymbol{\Sigma}_{2}^{\mathbf{x}},\boldsymbol{\Sigma}_{1}^{\mathbf{y}},\boldsymbol{\Sigma}_{2}^{\mathbf{y}}$) are assumed to be nonanticipative (adapted) functionals over the measurable time-function cylinder
$(C^{0,k}_{T},\mathscr{B}^{k}_{T}):=\big{(}[0,T]\times C^{0}([0,T];\mathbb{C}^{k}),\mathcal{B}([0,T])\otimes\mathcal{B}(C^{0}([0,T];\mathbb{C}^{k}))\big{)},$
@@
$\mathscr{B}^{k}_{T}=\mathcal{B}([0,T])\otimes\mathcal{B}(C^{0}([0,T];\mathbb{C}^{k}))=\sigma\left(\big{\{}A\times B:A\in\mathcal{B}([0,T]),B\in\mathcal{B}(C^{0}([0,T];\mathbb{C}^{k}))\big{\}}\right),$
@@
with $\mathcal{B}(C^{0}([0,T];\mathbb{C}^{k}))$ being the $\sigma$-algebra generated by the topology of compact convergence on the space of continuous functions from $[0,T]$ to $\mathbb{C}^{k}$, $C^{0}([0,T];\mathbb{C}^{k})$. This means that if $g(t,\mathbf{x})$ denotes a standard element of these functionals, for each $t\in[0,T]$, it is $(\mathcal{F}_{t}^{\mathbf{x}};\mathbb{C})$-measurable. It is important to emphasize here the fact that in a CGNS, the unobservable component $\mathbf{y}$ enters the dynamics in a conditionally linear manner, whereas the observable process $\mathbf{x}$ can enter into the coefficients of both equations in any $\mathcal{F}^{\mathbf{x}}_{t}$-measurable way.
@@
Many CTNDSs fit into the CGNS modelling framework. Some well-known classes of these systems are physics-constrained nonlinear stochastic models (for example the noisy versions of the Lorenz models, low-order models of Charney-DeVore flows, and a paradigm model for topographic mean flow interaction) [51, 52], stochastically coupled reaction-diffusion models used in neuroscience and ecology (for example the stochastically coupled FitzHugh-Nagumo models and the stochastically coupled SIR epidemic models), and spatiotemporally multiscale models for turbulence, fluids, and geophysical flows (for example the Boussinesq equations with noise and the stochastically forced rotating shallow-water equation) [38]. This modelling framework has also been exploited to develop realistic low-order stochastic models for the Madden-Julian oscillation (MJO) and Arctic sea ice [41, 53].
@@
In addition to modeling many natural phenomena, the CGNS framework and its closed analytic DA formulae have been applied to study many theoretical and practical problems. It has been utilized to develop a nonlinear LDA algorithm, allowing rigorous analysis to study model error and uncertainty, as well as recovery of turbulent ocean flows with noisy observations from Lagrangian tracers [42, 47, 48, 54]. The analytically solvable DA scheme has also been applied to the prediction and state estimation of the non-Gaussian intermittent time series of the MJO and the monsoon intraseasonal variabilities, in addition to the filtering of the stochastic skeleton model for the MJO [37, 41, 55, 56, 57, 58]. Notably, the efficient DA procedure also helps develop a rapid algorithm to solve the high-dimensional Fokker-Planck equation [59, 60]. Worth highlighting is that the ideas of the CGNS modelling framework and the associated DA procedures have also been applied to develop cheap exactly-solvable forecast models in dynamic stochastic superresolution of sparsely observed turbulent systems [61, 62], build stochastic superparameterization for geophysical turbulence [63, 64, 65, 66], and design efficient multiscale DA schemes via blended particle filters for high-dimensional chaotic systems [67].
@@
A set of sufficient regularity conditions needs to be assumed a-priori so that the main result of the CGNS framework can be established. We enforce the same set of assumptions as in the work of Andreou & Chen [68], which we outline in Appendix A, with the rationale behind the adoption of each one being provided there (and in the references therein). These conditions are then sufficient enough to show that the posterior distribution of the unobserved variables, when conditioning on the observational data, is Gaussian, as stated in the following theorem; this is exactly why the system in (2.1)–(2.2) is called a CGNS of equations. For the following, we abuse notation for clarity and instead write $\big{(}\boldsymbol{\cdot}\big{|}\mathbf{x}(s),s\leq t\big{)}$ to indicate the fact that we are conditioning on the observable $\sigma$-algebra at time $t$, $\big{(}\boldsymbol{\cdot}\big{|}\mathcal{F}_{t}^{\mathbf{x}}\big{)}$.
@@
Let $(\mathbf{x}(t),\mathbf{y}(t))$ satisfy (2.1)–(2.2) and assume that the regularity conditions (1)(6) found in Appendix A hold. Additionally, assume that the initial conditional distribution $\mathbb{P}\big{(}\mathbf{y}(0)\leq\boldsymbol{\alpha}_{0}\big{|}\mathbf{x}(0)\big{)}$ [^1] The event $\{\mathbf{y}(s)\leq\boldsymbol{\alpha}=(\alpha_{1},\ldots,\alpha_{l})^{\mathtt{T}}\}$ is to be understood coordinate-wise, i.e., $\{y_{1}(s)\leq\alpha_{1},\ldots,y_{l}(s)\leq\alpha_{l}\}$, with the linear partial ordering assumed on $\mathbb{C}$, i.e., $y_{j}(s)\leq\alpha_{j}\Leftrightarrow\big{(}\mathrm{Re}(y_{j}(s))<\mathrm{Re}(\alpha_{j})\big{)}\vee\big{(}\mathrm{Re}(y_{j}(s))=\mathrm{Re}(\alpha_{j})\wedge\mathrm{Im}(y_{j}(s))\leq\mathrm{Im}(\alpha_{j})\big{)}$. be ($\mathbb{P}$-a.s.) Gaussian, $\mathcal{N}_{l}(\boldsymbol{\mu}_{\text{f}}(0),\mathbf{R}_{\text{f}}(0))$, and mutually independent from the Wiener processes $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, where
$\boldsymbol{\mu}_{\text{f}}(0):=\mathbb{E}\left[\mathbf{y}(0)\big{|}\mathbf{x}(0)\right]\quad\text{and}\quad\mathbf{R}_{\text{f}}(0):=\mathbb{E}\left[(\mathbf{y}(0)-\boldsymbol{\mu}_{\text{f}}(0))(\mathbf{y}(0)-\boldsymbol{\mu}_{\text{f}}(0))^{\dagger}\big{|}\mathbf{x}(0)\right],$
with $\boldsymbol{\cdot}^{\dagger}$ denoting the Hermitian transpose operator. Furthermore, assume $\mathbb{P}\left(\mathrm{tr}(\mathbf{R}_{\text{f}}(0))<+\infty\right)=1$, where $\mathrm{tr}(\boldsymbol{\cdot})$ denotes the trace operator, meaning that the initial estimation mean-square error between $\mathbf{y}(0)$ and $\boldsymbol{\mu}_{\text{f}}(0)$ is almost surely finite. Then, for any $t_{j}$ such that $0\leq t_{1}<t_{2}<\cdots<t_{n}\leq t$, with $t\in[0,T]$, and $\boldsymbol{\alpha}_{1},\ldots,\boldsymbol{\alpha}_{n}\in\mathbb{C}^{l}$, the conditional distribution
@@
is ($\mathbb{P}$-a.s.) Gaussian.
@@
This is Theorem 12.6 in Liptser & Shiryaev [36], which is the multi-dimensional analog of Theorem 11.1. For the analogous result in the case of discrete time, see Theorem 13.3, with the respective sufficient assumptions given in Subchapter 13.2.1. Thorough details are also provided in Kolodziej [69] for the continuous-time case. The proof, regardless of continuous- or discrete-time, uses the conditional characteristic function method and a conditional version of the law-uniqueness theorem [69, 70]. ∎
@@ 2.2 Analytically Solvable Filter and Smoother Posterior Distributions
Let the assumptions of Theorem 2.1 and the additional regularity conditions (7)(9), which are outlined in Appendix A, to hold. Then for any $t\in[0,T]$ the $\mathcal{F}^{\mathbf{x}}_{t}$-measurable Gaussian statistics of the Gaussian conditional distribution
$\mathbb{P}\big{(}\mathbf{y}(t)\leq\boldsymbol{\alpha}\big{|}\mathbf{x}(s),s\leq t\big{)}\overset{{\,\rm d}}{\sim}\mathcal{N}_{l}(\boldsymbol{\mu}_{\text{{f}}}(t),\mathbf{R}_{\text{{f}}}(t)),$
where
$\boldsymbol{\mu}_{\text{{f}}}(t):=\mathbb{E}\left[\mathbf{y}(t)\big{|}\mathbf{x}(s),s\leq t\right]\quad\text{and}\quad\mathbf{R}_{\text{{f}}}(t):=\mathbb{E}\left[(\mathbf{y}(t)-\boldsymbol{\mu}_{\text{{f}}}(t))(\mathbf{y}(t)-\boldsymbol{\mu}_{\text{{f}}}(t))^{\dagger}\big{|}\mathbf{x}(s),s\leq t\right],$
@@
Furthermore, if the initial covariance matrix $\mathbf{R}_{\text{{f}}}(0)$ is positive-definite ($\mathbb{P}$-a.s.), then all the matrices $\mathbf{R}_{\text{{f}}}(t)$, for $t\in[0,T]$, remain positive-definite ($\mathbb{P}$-a.s.).
@@
This is Theorem 12.7 in Liptser & Shiryaev [36], which is the multi-dimensional analog of Theorems 12.1 and 12.3. For the analogous result in the case of discrete time, see Theorem 13.4, which outlines the corresponding optimal recursive nonlinear filter difference equations, with the respective sufficient assumptions given in Subchapter 13.2.1. Thorough details are also provided in Kolodziej [69] for the continuous-time case. For a martingale-free proof see Andreou & Chen [68] (although it requires the additional assumptions (10)(11) found in Appendix A, as to pass from the discrete-time filter to the continuous one via a formal limit). ∎
@@
The form of the filter mean equation can be intuitively explained within the Bayesian inference framework for DA. The first two terms on the right-hand side of (2.3), namely $\boldsymbol{\Lambda}^{\mathbf{y}}\boldsymbol{\mu}_{\text{f}}+\mathbf{f}^{\mathbf{y}}$, represent the forecast mean, derived from the process of the unobserved variables in (2.2). The remaining terms account for correcting the posterior mean state, incorporating information from partial observations. The matrix factor in front of the innovation, or pre-fit measurement residual ${\,\rm d}\mathbf{x}(t)-(\boldsymbol{\Lambda}^{\mathbf{x}}\boldsymbol{\mu}_{\text{{f}}}+\mathbf{f}^{\mathbf{x}}){\,\rm d}t$, is analogous to the Kalman gain in classical Kalman filter theory: ${(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}}+\mathbf{R}_{\text{{f}}}(\boldsymbol{\Lambda}^{\mathbf{x}})^{\dagger})(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{-1}}$. This factor determines the weight of the observations when updating the model-predicted state. Even in cases where the variables $\mathbf{y}$ of the unobserved process do not explicitly depend on $\mathbf{x}$, such as in the Lagrangian DA, the observational process (2.1) still couples the observed and unobserved components. This coupling allows the observations to impact state estimation and correct the forecast. Finally, observe that the equation driving the evolution of the covariance tensor $\mathbf{R}_{\text{f}}$ is a random Riccati equation [71, 72, 73], since the coefficients depend on the observable random variables $\mathbf{x}$.
@@
As already stated, the CGNS framework also enjoys closed analytic formulae for the recovery of the optimal smoother state estimation. The smoother posterior distribution at time $t\in[0,T]$ exploits the observational information in the entire period $[0,T]$ and therefore it allows for a more accurate, less uncertain, and unbiased estimated state compared to the filter solution. Solving the optimal nonlinear smoother equations requires applying a forward pass (filtering) from $t=0$ to $t=T$, which is then followed by a backward pass (smoothing) from $t=T$ to $t=0$ [15, 16, 17, 18, 19]. In the theorem that follows, the optimal nonlinear smoother equations, which run backwards in time, are showed, with the subscript “ s ” is denoting the smoother conditional Gaussian statistics, which appropriately stands for smoother. The smoother conditional Gaussian statistics are also known as the smoother posterior mean and smoother posterior covariance under the Bayesian inference dynamics framework.
@@
Let the assumptions of Theorem 2.1 and the additional regularity conditions (7)(9), which are outlined in Appendix A, to hold. In addition, assume

### v2

@@ 2.1 Conditionally Gaussian Processes
In its most general form, a conditional Gaussian system of processes consists of two diffusion-type processes defined by the following nonlinear system of stochastic differentials given in Itô form [33, 34]:
$\mathrm{d}\mathbf{x}(t)$ $=\big(\boldsymbol{\Lambda}^{\mathbf{x}}(t,\mathbf{x})\mathbf{y}(t)+\mathbf{f}^{\mathbf{x}}(t,\mathbf{x})\big)\mathrm{d}t+\boldsymbol{\Sigma}_{1}^{\mathbf{x}}(t,\mathbf{x})\mathrm{d}\mathbf{W}_{1}(t)+\boldsymbol{\Sigma}_{2}^{\mathbf{x}}(t,\mathbf{x})\mathrm{d}\mathbf{W}_{2}(t),$ (2.1)
$\mathrm{d}\mathbf{y}(t)$ $=\big(\boldsymbol{\Lambda}^{\mathbf{y}}(t,\mathbf{x})\mathbf{y}(t)+\mathbf{f}^{\mathbf{y}}(t,\mathbf{x})\big)\mathrm{d}t+\boldsymbol{\Sigma}_{1}^{\mathbf{y}}(t,\mathbf{x})\mathrm{d}\mathbf{W}_{1}(t)+\boldsymbol{\Sigma}_{2}^{\mathbf{y}}(t,\mathbf{x})\mathrm{d}\mathbf{W}_{2}(t),$ (2.2)
@@
are two mutually independent complex-valued Wiener processes (i.e., both their real and imaginary parts are mutually independent real-valued Wiener processes with standardized covariances) and almost every path of $\mathbf{x}$ and $\mathbf{y}$ is in $C^{0}\big([0,T];\mathbb{C}^{k}\big)$ and $C^{0}\big([0,T];\mathbb{C}^{l}\big)$, respectively. The elements of the vector- and matrix-valued functions of multiplicative factors ($\boldsymbol{\Lambda}^{\mathbf{x}},\boldsymbol{\Lambda}^{\mathbf{y}}$), forcings ($\mathbf{f}^{\mathbf{x}},\mathbf{f}^{\mathbf{y}}$), and noise feedbacks ($\boldsymbol{\Sigma}_{1}^{\mathbf{x}},\boldsymbol{\Sigma}_{2}^{\mathbf{x}},\boldsymbol{\Sigma}_{1}^{\mathbf{y}},\boldsymbol{\Sigma}_{2}^{\mathbf{y}}$) are assumed to be nonanticipative (adapted) functionals over the measurable time-function cylinder
$\big(C^{0,k}_{T},\mathscr{B}^{k}_{T}\big):=\left([0,T]\times C^{0}\big([0,T];\mathbb{C}^{k}\big),\mathcal{B}\big([0,T]\big)\otimes\mathcal{B}\big(C^{0}\big([0,T];\mathbb{C}^{k}\big)\big)\right),$
@@
$\mathscr{B}^{k}_{T}=\mathcal{B}\big([0,T]\big)\otimes\mathcal{B}\big(C^{0}\big([0,T];\mathbb{C}^{k}\big)\big)=\sigma\left(\big\{A\times B:A\in\mathcal{B}\big([0,T]\big),B\in\mathcal{B}\big(C^{0}\big([0,T];\mathbb{C}^{k}\big)\big)\big\}\right),$
@@
with $\mathcal{B}\big(C^{0}\big([0,T];\mathbb{C}^{k}\big)\big)$ being the $\sigma$-algebra generated by the topology of compact convergence on the space of continuous functions from $[0,T]$ to $\mathbb{C}^{k}$, $C^{0}\big([0,T];\mathbb{C}^{k}\big)$. It is important to emphasize here the fact that in a CGNS, the unobservable component $\mathbf{y}$ enters the dynamics in a conditionally linear manner, whereas the observable process $\mathbf{x}$ can enter into the coefficients of both equations in any measurably nonlinear way.
@@
Many CTNDSs fit into the CGNS modeling framework. Some well-known classes of these systems are physics-constrained nonlinear stochastic models (for example the noisy versions of the Lorenz models, low-order models of Charney–DeVore flows, and a paradigm model for topographic mean flow interaction) [45, 46], stochastically coupled reaction-diffusion models used in neuroscience and ecology (for example the stochastically coupled FitzHugh–Nagumo models and the stochastically coupled SIR epidemic models), and spatiotemporally multiscale models for turbulence, fluids, and geophysical flows (for example the Boussinesq equations with noise and the stochastically forced rotating shallow-water equation) [34]. This modeling framework has also been exploited to develop realistic low-order stochastic models for the Madden–Julian oscillation (MJO) and Arctic sea ice [37, 47].
@@
In addition to modeling many natural phenomena, the CGNS framework and its closed analytic DA formulae have been applied to study many theoretical and practical problems. It has been utilized to develop a nonlinear Lagrangian data assimilation algorithm, allowing rigorous analysis to study model error and uncertainty, as well as recovery of turbulent ocean flows with noisy observations from Lagrangian tracers [41, 42, 43]. The analytically solvable DA scheme has also been applied to the prediction and state estimation of the non-Gaussian intermittent time series of the MJO and the monsoon intraseasonal variabilities, in addition to the filtering of the stochastic skeleton model for the MJO [37, 48, 44]. Notably, the efficient DA procedure also helps develop a rapid algorithm to solve the high-dimensional Fokker–Planck equation [49, 50]. Worth highlighting is that the ideas of the CGNS modeling framework and the associated DA procedures have also been applied to develop cheap exactly-solvable forecast models in dynamic stochastic superresolution of sparsely observed turbulent systems [51, 52], build stochastic superparameterization for geophysical turbulence [53], and design efficient multiscale DA schemes via blended particle filters for high-dimensional chaotic systems [54].
@@
A set of sufficient regularity conditions needs to be assumed a-priori so that the main results of the CGNS framework can be established. We enforce the same set of assumptions as in the work of Andreou & Chen [55], which we outline in Appendix A, with the rationale behind the adoption of each one being provided there (and in the references therein). These conditions are sufficient to show that the posterior distributions of the unobserved variables, when conditioning on the observational data, are Gaussian, as stated in the following theorem; this is exactly why (2.1)–(2.2) is called a CGNS. For the following, we abuse notation for clarity and instead write $\big(\cdot\big|\mathbf{x}(s),s\leq t\big)$ to indicate the fact that we are conditioning on the observable $\sigma$-algebra at time $t$, $\big(\cdot\big|\mathcal{F}_{t}^{\mathbf{x}}\big)$.
@@
Let $\big(\mathbf{x}(t),\mathbf{y}(t)\big)$ satisfy (2.1)–(2.2) and assume that the regularity conditions (1)(8) in Appendix A hold. Additionally, assume that the initial conditional distribution $\mathbb{P}\big(\mathbf{y}(0)\leq\boldsymbol{\alpha}_{0}\big|\mathbf{x}(0)\big)$ [^1] The event $\big\{\mathbf{y}(s)\leq\boldsymbol{\alpha}=(\alpha_{1},\ldots,\alpha_{l})^{\mathtt{T}}\big\}$ is to be understood coordinate-wise: $\big\{j=1,\ldots,l:\big(\mathrm{Re}(y_{j}(s))\leq\mathrm{Re}(\alpha_{j}),\mathrm{Im}(y_{j}(s))\leq\mathrm{Im}(\alpha_{j})\big)\big\}$. is ($\mathbb{P}$-almost surely) Gaussian, $\mathcal{N}_{l}\big(\boldsymbol{\mu}_{\text{f}}(0),\mathbf{R}_{\text{f}}(0)\big)$ [^2] Whenever we refer to a complex-valued multivariate Gaussian or normal distribution in this work, we mean a circularly-symmetric one (i.e., with a zero pseudo-covariance or relation matrix) [56]., and mutually independent from the Wiener processes $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, where
$\boldsymbol{\mu}_{\text{f}}(0):=\mathbb{E}\big[\mathbf{y}(0)\big|\mathbf{x}(0)\big]\quad\text{and}\quad\mathbf{R}_{\text{f}}(0):=\mathbb{E}\big[(\mathbf{y}(0)-\boldsymbol{\mu}_{\text{f}}(0))(\mathbf{y}(0)-\boldsymbol{\mu}_{\text{f}}(0))^{\dagger}\big|\mathbf{x}(0)\big],$
with $\cdot^{\dagger}$ denoting the Hermitian transpose operator. Furthermore, assume $\mathbb{P}\big(\mathrm{tr}(\mathbf{R}_{\text{f}}(0))<+\infty\big)=1$, where $\mathrm{tr}(\cdot)$ denotes the trace operator, meaning that the initial estimation mean-square error between $\mathbf{y}(0)$ and $\boldsymbol{\mu}_{\text{f}}(0)$ is almost surely finite. Then, for any $t_{j}$ such that $0\leq t_{1}<t_{2}<\cdots<t_{n}\leq t$, with $t\in[0,T]$, and $\boldsymbol{\alpha}_{1},\ldots,\boldsymbol{\alpha}_{n}\in\mathbb{C}^{l}$, the conditional distribution
@@
is ($\mathbb{P}$-almost surely) Gaussian.
@@
This is Theorem 12.6 in Liptser & Shiryaev [33], which is the multi-dimensional analog of Theorem 11.1. Thorough details are also provided in Kolodziej [57]. For the analogous result in the case of discrete time (i.e., where the CGNS consists of stochastic difference equations and $\mathbf{x}$ is observed at discrete times instants), see Theorem 13.3 of Liptser & Shiryaev [33], with the respective sufficient assumptions given in Subchapter 13.2.1.
The proof, regardless of continuous- or discrete-time, uses the conditional characteristic function method and a conditional version of the law-uniqueness theorem [57, 58]. ∎
@@ 2.2 Analytically Solvable Filter and Smoother Posterior Distributions
Let the assumptions of Theorem 2.1 and the additional regularity conditions (9), (11), and (12), which are outlined in Appendix A, to hold. Then, for any $t\in[0,T]$, the $\mathcal{F}^{\mathbf{x}}_{t}$-measurable Gaussian statistics of the Gaussian conditional distribution
$\mathbb{P}\big(\mathbf{y}(t)\leq\boldsymbol{\alpha}\big|\mathbf{x}(s),s\leq t\big)\overset{{\,\rm d}}{\sim}\mathcal{N}_{l}\big(\boldsymbol{\mu}_{\text{{f}}}(t),\mathbf{R}_{\text{{f}}}(t)\big),$
defined as
$\boldsymbol{\mu}_{\text{{f}}}(t):=\mathbb{E}\big[\mathbf{y}(t)\big|\mathbf{x}(s),s\leq t\big]\quad\text{and}\quad\mathbf{R}_{\text{{f}}}(t):=\mathbb{E}\big[(\mathbf{y}(t)-\boldsymbol{\mu}_{\text{{f}}}(t))(\mathbf{y}(t)-\boldsymbol{\mu}_{\text{{f}}}(t))^{\dagger}\big|\mathbf{x}(s),s\leq t\big],$
@@
Furthermore, if the initial covariance matrix $\mathbf{R}_{\text{{f}}}(0)$ is positive-definite ($\mathbb{P}$-almost surely), then all the matrices $\mathbf{R}_{\text{{f}}}(t)$, for $t\in[0,T]$, remain positive-definite ($\mathbb{P}$-almost surely).
@@
This is Theorem 12.7 in Liptser & Shiryaev [33], which is the multi-dimensional analog of Theorems 12.1 and 12.3. Thorough details are also provided in Kolodziej [57]. For the analogous result in the case of discrete time, see Theorem 13.4 of Liptser & Shiryaev [33], which outlines the corresponding optimal recursive nonlinear filter difference equations, with the respective sufficient assumptions given in Subchapter 13.2.1. For a martingale-free proof see Theorem 2 in Andreou & Chen [55] (although it additionally requires assumption (10) found in Appendix A, as to pass from the discrete-time filter to the continuous one via a formal limit). ∎
@@
The form of the filter mean equation can be intuitively explained within the Bayesian inference framework for DA. The first two terms on the right-hand side of (2.3), namely $\boldsymbol{\Lambda}^{\mathbf{y}}\boldsymbol{\mu}_{\text{f}}+\mathbf{f}^{\mathbf{y}}$, represent the forecast mean, derived from the process of the unobserved variables in (2.2). The remaining terms account for correcting the prior mean state, incorporating information from partial observations. The matrix factor in front of the innovation, or pre-fit measurement residual ${\,\rm d}\mathbf{x}(t)-(\boldsymbol{\Lambda}^{\mathbf{x}}\boldsymbol{\mu}_{\text{{f}}}+\mathbf{f}^{\mathbf{x}}){\,\rm d}t$, is analogous to the Kalman gain in classical Kalman filter theory: $(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}}+\mathbf{R}_{\text{{f}}}(\boldsymbol{\Lambda}^{\mathbf{x}})^{\dagger})(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{-1}$}. This factor determines the weight of the observations when updating the model-predicted state. Even in cases where the variables $\mathbf{y}$ of the unobserved process do not explicitly depend on $\mathbf{x}$, such as in the Lagrangian DA, the observational process (2.1) still couples the observed and unobserved components. This coupling allows the observations to impact state estimation and correct the forecast. Finally, observe that the equation driving the evolution of the covariance tensor $\mathbf{R}_{\text{f}}$ is a random Riccati equation [59, 60], since the coefficients depend on the observable random variables $\mathbf{x}$.
@@
As already stated, the CGNS framework also enjoys closed analytic formulae for the recovery of the optimal smoother state estimation. The smoother posterior distribution at time $t\in[0,T]$ exploits the observational information in the entire period $[0,T]$ and therefore it allows for a more accurate and less uncertain estimated state compared to the filter solution. Solving the optimal nonlinear smoother equations requires applying a forward pass (filtering) from $t=0$ to $t=T$, which is then followed by a backward pass (smoothing) from $t=T$ to $t=0$ [14, 15, 16, 17]. In the theorem that follows, the optimal nonlinear smoother equations, which run backwards in time, are showed, with the subscript “ s ” denoting the smoother conditional Gaussian statistics, which appropriately stands for smoother. The smoother conditional Gaussian statistics are also known as the smoother posterior mean and smoother posterior covariance under the Bayesian inference dynamics framework.
@@
Let the assumptions of Theorem 2.1 and the additional regularity conditions (9), (11), and (12), which are outlined in Appendix A, to hold. In addition, assume